I'd appriciate some help on this question sense i have trouble graphing.
QUESTION: graph each equation by writing two linear equations.
PROBLEM: y=|x-3| ??

Answer by Edwin McCravy(20086)

y = |x-3|

Find the vertex (turning point) by setting what is between
the absolute value bars = 0,    x-3 = 0
                                  x = 3

The two equations are

y = x-3 for x ≧ 3    and y = -(x-3) or y = -x+3 for x < 3

Get the vertex and a point on each side of the vertex:

x | y
2 | 1
3 | 0
4 | 1



Edwin